OpenStudy (anonymous):

use implicit differentiation to find the equation of the tangent line in the form y=mx+b at the point (-2,3) x^2+y^2=13

OpenStudy (anonymous):

i worked out the whole thing and i came up with y=2/3x+13 but i really don't think i knew how to find my b, but I'm not sure that was my only problem

OpenStudy (anonymous):

Alright to solve this take the derivative 2x + y*y' = 0 = y' = y/2x now to find the equation of the tangent line we use the formula f'(x)(x-a)+f(x) we have the point (-2,3) so we know f(-2) = 3 and the x value we are looking at is -2 so a = -2 so we have f'(x)(x+2) + 3 now we need to find f'(-2) and we have our equation of the tangent line to do this sub in the point into the equation and solve for y' y' = y/2x y' = 3/2(-2) y' = -3/4 we now have f'(-2) = -3/4 so the equation is -3/4(x+2) + 3 = y

OpenStudy (anonymous):

Sorry I mean (-3/4)(x+2) + 3 = y

OpenStudy (anonymous):

the equation I provided is the same as the equation of the line, I have proof but I'm far too lazy to show it.

OpenStudy (anonymous):

Also it is y'y becuase I applied chain rule

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

here is an example of implicit differentiation